OpenStudy (anonymous):

Kyle has 22 more cards than 3 times the number of cards that Stan owns. Together, they own less than 58 cards. How many cards are in Stan's collection? Stan owns 9 cards or less. Stan owns 9 cards or more. Stan owns less than 9 cards. Stan owns more than 9 cards.

undeadknight26 (undeadknight26):

For this we can just plug and play. For example lets take A: 9 x 3 = 27 + 22 = 49 + 9 = 58. I'd say A. That was a fail of an example.

OpenStudy (phi):

Kyle has 22 more cards than let K stand for the number of cards Kyle has let X = the number Stan has 20 more than means: 22+ 3 times the number of cards that Stan owns. means 3 times X or 3x so K= 22+3x

OpenStudy (phi):

Together, they own less than 58 cards. that means we add them up K+X < 58 because we know K = 22+3X we can also write it as 22+3X+X < 58 or, add 3 X's and 1 X to get 4 X's 22+4X < 58 now add -22 to both sides 22-22 +4X < 58 - 22 4X< 36